in a match the column question, column 1 conatains 10 questions and column 2 contains 10 answers written in some arbitrary order. in how many ways a student can answer this question so that exactly 6 of his matchings are correct?

Grade:

2 Answers

Sher Mohammad IIT Delhi
10 years ago
4 wrong question can be selected in 10C4 ways and this can be permuted in 4! ways,
total ways 10C4*4!


sher mohammad
b.tech, iit delhi
sneha parmar
3 years ago
we know the formula of dearrangment we have to use it
acc to ques. 6 ans are matching so we will select 6 out of 10 that is 10c6rest we will arrange in 4! ways after putting in formula we would get
=10c6 * 4!( 1/2! -1/3! +1/4!)
=210 *4! (12-4+1/4!)
(4! and 4! will get cabcel remaining will be)
=210(12-3)
=210*9
=1890 (answer)
 
by Sneha Parmar ...to be iitian